### v2

# Prometheus Mind: Retrofitting Memory to Frozen Language Models

Abstract
Adding memory to pretrained language models typically requires architectural changes or weight modification. We present Prometheus Mind, which retrofits memory to a frozen Qwen3-4B using 11 modular adapters (530MB, 7% overhead)—fully reversible by removing the adapters.
Building this system required solving four problems: (1) Extraction—we develop Contrastive Direction Discovery (CDD), which finds semantic directions via minimal pairs without labeled data. (2) Training—end-to-end optimization collapses; stage-wise training of each adapter on simple proxy tasks succeeds. (3) Injection—learned encoders fail to generalize; we find that lm_head.weight rows already provide the mapping we need, requiring no training. (4) Hidden state collapse—transformers make “wife” and “brother” 0.98+ similar; we train projections to recover distinction (0.98 $\rightarrow$ 0.09).
On PrometheusExtract-132 (132 cases spanning 10 difficulty categories), the system achieves 94.4% retrieval on clean inputs (n=54, 95% CI: [84.9%, 98.1%]), 57.6% overall, and 19.4% on adversarial inputs with ellipsis, filler words, or implicit subjects (n=36). The primary bottleneck is relation classification (47.3% accuracy), responsible for most extraction errors.
1 Introduction
Large language models exhibit remarkable capabilities but fundamentally lack persistent memory. Their only form of “memory” is the context window—a fixed-length buffer that is cleared between sessions and cannot accumulate knowledge over time. A user who tells the model their name in one conversation must repeat it in the next; facts learned in morning interactions are forgotten by afternoon. This limitation motivates a growing body of work on memory-augmented language models.
Current approaches to adding memory fall into two categories, each with significant limitations. Retrieval-Augmented Generation (RAG) systems retrieve relevant text from external databases and inject it into the context window as additional tokens [7]. While effective, this approach treats memory as input—the model must “read” the retrieved text through its normal attention mechanism, competing with the user’s actual query for attention budget. The model does not think with the memory; it reads it. In contrast, architectural approaches like Titans [1] integrate neural memory modules directly into the transformer architecture, enabling surprise-driven memorization at test time. However, these approaches require modifying the base model architecture or weights, making them non-reversible and incompatible with existing pretrained checkpoints.
We propose a third path: memory as a removable module for frozen language models. Our system, Prometheus Mind, adds memory capabilities to a frozen Qwen3-4B model through 11 small adapter networks (530MB total, 7% parameter overhead) and hidden state injection hooks. Critically, unlike RAG which injects memory as context tokens, we inject memory directly into the model’s attention mechanism at layer 35 of 36 (97% depth)—the model attends to the memory through its native attention computation rather than reading it as text. And unlike Titans, our approach requires no architectural changes: the base model weights remain frozen, and removing the adapters restores the original model exactly.
A key insight from our development is that simply training an adapter to inject memory is insufficient—the frozen model learns to ignore injected signals. We initially explored LoRA adapters on attention layers to create a “reception pathway,” achieving 87.5% accuracy. However, our final breakthrough was simpler: by using the model’s own lm_head.weight rows as value vectors (Identity V), the frozen model generates correct outputs without modifying base model weights. Unlike learned value encoders that memorize training examples and fail on unseen facts, Identity V generalizes to any token in the vocabulary because it uses the model’s native output representations.
The journey to memory.
Retrofitting memory to a frozen LLM requires solving four problems in sequence. Each solution became a contribution:
  1. 1.
    Extracting facts: How do we identify what to store without labeled training data? We develop Contrastive Direction Discovery (CDD), a self-supervised method that discovers semantic directions by contrasting minimal pairs for subject, verb, object, and temporal extraction.
  2. 2.
    Training reliably: Joint optimization of 11 adapters collapses to random performance. We find that stage-wise training—training each adapter in isolation on a simple proxy task—produces a working system while end-to-end training fails.
  3. 3.
    Injecting answers: How do we inject retrieved facts such that the model generates appropriate responses? Learned value encoders memorize training data (100% train, 0% test). We discover that lm_head.weight rows are already optimal value vectors, requiring no training and generalizing to any vocabulary token.
  4. 4.
    Fixing hidden state collapse: Transformer hidden states collapse semantically distinct concepts (“wife” vs “brother” have 0.98+ cosine similarity). We train projections using embed_tokens as supervision, recovering distinction (0.98 $\rightarrow$ 0.09).
The result: memory for a frozen Qwen3-4B model via 11 modular adapters (530MB, 7% overhead), achieving 94.4% retrieval on clean inputs (n=54, 95% CI: [84.9%, 98.1%]), 57.6% overall, and 19.4% on adversarial inputs (n=36). We evaluate on PrometheusExtract-132, a benchmark of 132 cases spanning 10 difficulty categories with 95% confidence intervals for all metrics.
Limitations acknowledged upfront.
Relation classification remains the primary bottleneck (47.3% accuracy), accounting for 89% of extraction errors. Performance on complex inputs (multi-subject, multi-object) is poor (0% L6 accuracy). These limitations are analyzed in Section 6.
2 Related Work
2.1 Memory-Augmented Language Models
Recent work has explored various approaches to adding memory capabilities to language models. Titans [1] introduces a neural long-term memory module that learns to memorize at test time using a “surprise-driven” mechanism—events that violate expectations are stored more strongly. While Titans achieves impressive results scaling to 2M+ token contexts, it requires architectural modifications to the transformer, making it incompatible with existing pretrained models. LongMem [15] takes a different approach, using a frozen backbone as a memory encoder with an adaptive residual side-network for retrieval. This preserves the base model but still requires training a substantial side network. Mem0 [3] provides a scalable memory layer via external API. [^1] Mem0 reports 26% improvement on their multi-session dialogue benchmark; we cite their numbers without independent verification. However, Mem0 operates through context window injection rather than hidden state modification. LAVO [17] demonstrates that projecting keys into orthogonal slot space enables collision-free memory storage in linear attention. Our hierarchical memory structure builds on this principle.
Our approach differs fundamentally: we retrofit memory to a completely frozen model using only small adapter networks, with memory injected directly into hidden states rather than the context window. Table 1 summarizes these differences.
Table 1: Comparison of memory-augmented LLM approaches.
System Base Model Injection Point Reversible Training
Titans Modified Architectural No End-to-end
Mem0 Unchanged Context window Yes API-based
RAG Unchanged Context window Yes Retriever only
LongMem Frozen Residual network Partial Side network
Prometheus Frozen Hidden states Yes Stage-wise
2.2 Retrieval-Augmented Generation
RAG systems [7, 5] address knowledge limitations by retrieving relevant documents and prepending them to the context. Recent surveys  [4] categorize RAG architectures into retriever-centric, generator-centric, and hybrid designs. While RAG effectively grounds generation in external knowledge, it fundamentally differs from our approach in where memory is injected:
RAG: Memory $\rightarrow$ Tokens $\rightarrow$ Attention $\rightarrow$ Output
Ours: Memory $\rightarrow$ Hidden States $\rightarrow$ Layers $\rightarrow$ Output
In RAG, the model must allocate attention to reading the retrieved text. In our approach, memory directly modifies the model’s internal representations—the model thinks with the memory rather than reading it.
2.3 Adapter-Based Methods
Low-Rank Adaptation (LoRA) [6] enables efficient fine-tuning by freezing pretrained weights and adding small trainable low-rank matrices. This approach has been extended to serving multiple adapters concurrently [11] and continual learning settings. We adopt a similar philosophy—keeping the base model frozen while adding modular components—but apply it to memory rather than task adaptation. Our adapters perform extraction, routing, and retrieval rather than task-specific fine-tuning.
2.4 Activation Engineering and Steering Vectors
Activation Addition (ActAdd) [14] demonstrates that model behavior can be controlled by adding “steering vectors” computed from contrastive prompt pairs. Subsequent work has refined this with conceptors [9] (ellipsoidal regions rather than single vectors) and feature-guided approaches  [12] operating in sparse autoencoder latent spaces. Our Contrastive Direction Discovery (CDD) method shares the core insight that meaningful directions can be found by contrasting activations, but applies it to extraction rather than steering: we discover directions that isolate semantic roles (subject, verb, object) rather than behaviors.
2.5 Linear Probing and Semantic Subspaces
Work on probing transformer representations [2] has shown that semantic information is encoded in approximately linear subspaces  [13, 8]. Recent analysis reveals that transformers exhibit near-perfect linear relationships between sequential layers [10]. Our CDD method builds on these insights, using minimal pairs to discover linear directions that separate semantic roles. Additionally, our finding that hidden states “collapse” semantically distinct concepts relates to work on length-induced embedding collapse [18], though we observe collapse across semantic categories rather than sequence lengths.
3 Background
3.1 Transformer Language Models
A transformer language model processes input tokens $x_{1},\ldots,x_{n}$ through an embedding layer and $L$ transformer blocks. Each block applies multi-head self-attention followed by a feed-forward network, producing hidden states $\mathbf{h}_{i}^{(\ell)}\in\mathbb{R}^{d}$ at each layer $\ell$ and position $i$. The final hidden states are projected through lm_head, a linear layer $\mathbf{W}_{\text{lm}}\in\mathbb{R}^{V\times d}$ that produces logits over the vocabulary $V$.
Two components are particularly relevant to our approach:
  • embed_tokens: The input embedding matrix $\mathbf{W}_{\text{emb}}\in\mathbb{R}^{V\times d}$ maps token IDs to dense vectors. Importantly, this layer preserves token-level semantic distinctions that are later abstracted away in deeper layers.
  • lm_head: The output projection $\mathbf{W}_{\text{lm}}$ maps hidden states to vocabulary logits. Row $i$ of this matrix represents the “direction” that produces token ii—a key insight we exploit for memory injection.
3.2 Parameter-Efficient Adaptation
Parameter-efficient fine-tuning methods like LoRA [6] enable modifying pretrained models without full retraining. We initially explored LoRA to train the model to “accept” memory injections—adding low-rank updates to attention layers to create a reception pathway.
However, as described in Section 4.4, our Identity V discovery eliminates this need entirely: by using lm_head.weight rows directly as value vectors, the frozen model generates correct outputs without any weight modification. Our final system uses only small adapter networks for extraction and routing, keeping the base model completely frozen.
4 Method
This section presents our solutions to the four problems outlined in the introduction: extraction (Section 4.2), training (Section 4.3), injection (Section 4.4), and hidden state collapse (Section 4.5). We begin with a system overview.
4.1 System Overview
Prometheus Mind consists of three main components operating on a frozen Qwen3-4B model (Figure 1):
  1. 1.
    Extraction: Contrastive Direction Discovery (CDD) extracts semantic structure (subject, verb, object, temporal) from input text without labeled training data. Extracted facts are stored in a structured memory.
  2. 2.
    Retrieval: When a query arrives, adapter networks detect the query type and relation, then perform multi-hop reasoning over stored facts to retrieve the answer.
  3. 3.
    Injection: The retrieved answer is injected into the model’s the attention mechanism at layer 35 (97% depth) using lm_head.weight as the value encoder, enabling the model to generate responses grounded in memory.
The entire system uses 11 small adapter networks totaling 530MB (7% of the 8GB base model). All adapters can be enabled or disabled at runtime without affecting the frozen base model weights.
Memory Structure.
Facts are stored in a hierarchical slot structure inspired by LAVO [17]:
$\text{memory}[\text{type}][\text{slot}][\text{relation}]=\text{value\_embedding}$ (1)
where type $\in$ {people, places, organizations, concepts} (4 types), slot indexes entities within each type (64 slots per type), and relation $\in$ {profession, city, works_with, …} (13 relations). This yields 256 entity addresses (4 $\times$ 64) with 13 relation slots each.
[…]
Figure 1: Prometheus Mind architecture. Input flows through a frozen Qwen3-4B (left). CDD extracts facts into structured memory (right). At query time, multi-hop reasoning retrieves answers and injects them at layer 35 via attention K-V pairs, using Identity V (lm_head.weight).
4.2 Contrastive Direction Discovery (CDD)
CDD is a self-supervised method for extracting semantic structure from transformer hidden states without labeled training data. The key insight is that semantic distinctions are encoded in linear subspaces of the hidden state space.
4.2.1 The Minimal Pair Principle
When two sentences differ in exactly one semantic aspect, their difference vector isolates that aspect:
$\mathbf{d}=\mathbf{h}(\text{``Alice saw Bob''})-\mathbf{h}(\text{``Alice saw Carol''})$ (2)
The shared structure (“Alice saw [someone]”) cancels out, leaving only the object distinction (Bob vs Carol). By averaging many such difference vectors, we obtain a robust direction that points from one semantic role to another.
4.2.2 Hidden State CDD for Subject Extraction
For subject extraction, we combine two orthogonal directions:
Subject-Object Direction (Relational).
We contrast hidden states of tokens appearing before vs after the verb:
$\mathbf{d}_{\text{SO}}=\mathbb{E}[\mathbf{h}_{\text{subj}}-\mathbf{h}_{\text{verb}}]-\mathbb{E}[\mathbf{h}_{\text{obj}}-\mathbf{h}_{\text{verb}}]$ (3)
This separates tokens by their position relative to the verb.
Head-Modifier Direction (Absolute).
We contrast nouns vs determiners within the same phrase:
$\mathbf{d}_{\text{HM}}=\mathbb{E}[\mathbf{h}_{\text{noun}}]-\mathbb{E}[\mathbf{h}_{\text{det}}]$ (4)
This filters out determiners (“the”, “a”) that score high on the S-O direction but are not subjects.
Combined Scoring.
For each token $i$, we compute:
$\text{score}_{i}=\text{norm}(\mathbf{h}_{i}\cdot\mathbf{d}_{\text{SO}})+\text{norm}(\mathbf{h}_{i}\cdot\mathbf{d}_{\text{HM}})$ (5)
The token with the highest score is selected as the subject. On a small validation set (n=25), this two-direction approach achieves 92% accuracy (23/25), compared to 78% with the S-O direction alone. Note: These numbers are from development validation, not the main benchmark; sample size is insufficient for strong claims.
4.2.3 Attention CDD for Verb/Temporal Detection
Verbs present a challenge: they are defined by relationships rather than intrinsic properties, causing hidden state CDD to fail. We instead discover attention heads that specialize in attending to specific syntactic roles.
Head Discovery.
For each attention head $(l,h)$, we compute the difference in incoming attention between verbs and non-verbs:
$\Delta_{l,h}=\mathbb{E}[\text{attn}_{\rightarrow\text{verb}}]-\mathbb{E}[\text{attn}_{\rightarrow\text{non-verb}}]$ (6)
Heads with $\Delta>0$ attend preferentially to verbs. In Qwen3-4B, we identify five verb-selective heads: L6H2, L6H29, L5H20, L4H27, L5H23.
Inference.
At inference, we sum incoming attention from verb-selective heads for each token and select the maximum (excluding determiners and prepositions).
The same procedure discovers temporal-selective heads (L1H1, L6H4, L0H21, L0H12, L6H11). Object extraction combines object-selective attention heads with the H-M direction.
Development validation accuracies (small samples, not statistically robust): verb 91.7% (n=24), object 97.9% (n=48), temporal 92% (n=13). These numbers motivated the approach but are not rigorous evaluations. See Section 5 for benchmark results.
4.3 Stage-Wise Adapter Training
A critical finding from our development is that stage-wise training outperforms end-to-end training for modular memory systems. When we attempted joint optimization of all 11 adapters, training collapsed to near-random performance (roughly 1/11 accuracy for 11 relation types). Training each adapter in isolation on simple proxy tasks (50–200 examples per adapter) produced a working system.
Note: This is an observation from iterative development, not a controlled ablation. End-to-end training was abandoned early due to consistent failure. See Section 5 for benchmark results on the final system.
4.3.1 Why End-to-End Training Fails
End-to-end training of memory systems faces compounding difficulties:
  • Credit assignment: When the final output is wrong, it is unclear which component failed—extraction, storage, retrieval, or injection.
  • Gradient interference: Updates that improve one component may degrade another, leading to oscillation rather than convergence.
  • Sparse supervision: Most memory components receive gradients only when the final answer is evaluated, providing weak learning signal.
4.3.2 The Stage-Wise Approach
Instead of joint training, we train each adapter in isolation on a simple task with direct supervision:
  1. 1.
    Define the minimal task: What is the single decision this adapter must make? (e.g., “Is this word an entity?”)
  2. 2.
    Create targeted training data: 50–200 examples with clear ground truth for that specific decision.
  3. 3.
    Train to convergence: Simple cross-entropy loss, typically converging in 2–5 minutes on a single GPU.
  4. 4.
    Freeze and compose: Once trained, freeze the adapter and integrate it into the pipeline.
This approach treats the memory system as a composition of specialists rather than a monolithic network.
4.3.3 Adapter Inventory
Table 2 lists our 11 adapters grouped by function. Each adapter is a small network (typically 2–3 linear layers) trained on 50–200 examples.
Table 2: Adapter inventory. All adapters are trained stage-wise on small datasets.
Adapter Function Train Size Acc.
Extraction
BetaGate Entity vs function word 245 99%
IdentityAdapter Self-introduction detection 80 100%
UserAdapter First-person detection 120 98%
FirstPersonValueMask Multi-value extraction 150 100%
PairRelationClassifier Relation classification (17) 200 100%
Retrieval
IntermediateRelationDetector Query relation (17 classes) 916 99.7%
HopController Continue/stop decision 50 100%
AnswerSelector Return entity vs value 60 100%
QueryEntityExtractor Entity from question 80 100%
Routing
TypeClassifier Entity type (4 classes) 100 100%
RoutingAdapter Memory slot assignment 150 79%
MessageTypeClassifier Question/statement/chat 100 95%
Projections
UniversalProjection Fix word embedding collapse 500
SentenceProjection Fix sentence embedding collapse 683
4.4 Memory Injection via Attention K-V Pairs
Once a fact is retrieved from memory, it must be injected into the model’s processing such that the model generates an appropriate response. This seemingly simple problem proved surprisingly difficult.
4.4.1 Failed Approaches
We systematically explored multiple injection strategies before finding one that works:
  1. 1.
    Raw embedding injection: Averaging embeddings of the fact components and adding to hidden states. Result: Signal diluted to noise; model ignores injection entirely.
  2. 2.
    Trained context projector: Learning a projection from fact embeddings to hidden state space. Result: Projector converges but has no effect on generation—model learns to route around it.
  3. 3.
    Late-layer value injection: Directly replacing hidden states at layer 35/36 with fact embeddings. Result: Token repetition loops (“Mark Mark Mark Mark…”).
  4. 4.
    End-to-end trainable injector: Training the injection mechanism jointly with the model. Result: 100% accuracy on training data, 0% generalization to held-out facts.
The common failure mode is that the frozen model has learned to process its own hidden state distribution. Injected signals from external adapters lie outside this distribution and are ignored or cause instability.
4.4.2 The Attention K-V Breakthrough
The breakthrough came from reframing the problem: instead of asking “how do we inject a vector that makes the model output X?”, we asked “how do we let the model’s attention query our memory?”
The key insight is that the retrieval IS the injection—through the attention mechanism itself:
  • Q (query) carries “what is being asked” from the model to our adapters
  • V (value) carries “the answer” back from our adapters to the model
  • The model’s own attention mechanism handles the integration
We inject a memory key-value pair at layer 35 (97% depth) during the first generated token only. The model attends to this injected K-V pair through its normal attention computation, naturally incorporating the memory into its processing.
Why first-token-only?
Without this constraint, the model enters repetition loops (“Mark Mark Mark…”). By injecting only on the first generated token, the model outputs the injected value once, then continues naturally. This was a critical discovery during development.
4.4.3 Identity V: lm_head.weight as Memory Decoder
The final piece is the value encoder: how do we represent “the answer is chef” as a vector that, when attended to, causes the model to output “chef”?
We discovered that no training is needed for V. The model’s own lm_head.weight already contains the answer. (The K encoder—a small MLP mapping token embeddings to attention key space—does require training, but V is entirely training-free.)
$\mathbf{v}_{\text{memory}}=\alpha\cdot\frac{\mathbf{W}_{\text{lm}}[\text{token\_id}]}{\|\mathbf{W}_{\text{lm}}[\text{token\_id}]\|}\cdot\|\mathbf{h}\|$ (7)
where $\mathbf{W}_{\text{lm}}[\text{token\_id}]$ is row of lm_head corresponding to the answer token, and $\|\mathbf{h}\|$ matches the hidden state norm (critical: hidden states have norm $\sim$700 while lm_head rows have norm $\sim$1).
This works because lm_head was trained during pretraining to map hidden states to token probabilities—row $i$ is precisely the direction that produces token $i$. By using these rows as memory values, we achieve 100% generalization to any token in the vocabulary with zero V training.
[…]
4.5 Fixing Hidden State Collapse via Learned Projection
During development, we encountered a puzzling failure: the relation classifier could not distinguish “wife” from “brother” despite these being semantically distinct. Investigation revealed a fundamental problem with comparing transformer hidden states.
4.5.1 The Collapse Problem
Computing cosine similarity between hidden states of semantically distinct words yields surprisingly high values:
Word Pair Hidden State Sim. Expected
wife vs brother 0.99 Low
city vs profession 0.97 Low
Paris vs London 0.98 Medium
Yet the transformer knows these are different—it generates different continuations for “My wife is…” vs “My brother is…”. The distinction is preserved but hidden from naive cosine similarity.
4.5.2 SVD Analysis: Shared vs Distinctive Variance
SVD analysis of hidden states for relation words reveals the cause:
Components Variance Explained
Top 1 49.3%
Top 3 86.4%
Remaining 13.6%
86% of variance is shared structure—all these words are “relationship nouns,” and the transformer has learned this abstraction. The 14% distinctive information (which specific relationship) exists but is overwhelmed by the shared structure in cosine similarity.
Geometrically: all relation words point in roughly the same direction (the “relationship noun” direction), with small orthogonal components encoding which specific relation. Cosine similarity measures the dominant shared direction, not the distinctive components.
4.5.3 Learned Projection Solution
We solve this by learning a projection that amplifies distinctive information. The key insight is that embed_tokens (the input embedding layer) does preserve distinctions—it has not yet learned the abstractions that cause collapse.
We train a projection network $f_{\theta}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ to make hidden state similarities match embed_tokens similarities:
$\mathcal{L}=\text{MSE}\left(\text{sim}(f_{\theta}(\mathbf{h}_{i}),f_{\theta}(\mathbf{h}_{j})),\text{sim}(\mathbf{e}_{i},\mathbf{e}_{j})\right)$ (8)
where $\mathbf{h}_{i}$ are hidden states and $\mathbf{e}_{i}$ are embed_tokens embeddings.
The projection learns to extract the distinctive subspace, achieving:
Stage Avg Similarity (distinct words)
Before projection 0.981 (collapsed)
After projection 0.094 (distinct)
Target (embed_tokens) 0.094
We train two projections: UniversalProjection for word-level comparisons (relation matching), and SentenceProjection for question-level comparisons (query classification). The sentence projection uses relation labels as targets rather than embed_tokens, since questions about the same relation should cluster together regardless of specific words (“Where does my wife live?” should match “Where does my brother live?” for the “city” relation).
With projections applied, the relation classifier accuracy improves from 50% (random chance due to collapse) to 99.7%.
5 Experiments
5.1 Experimental Setup
Model.
We use Qwen3-4B [16] as our frozen base model (4B parameters, 36 layers, hidden dimension 2560). The model weights remain completely frozen throughout all experiments.
Adapters.
Our system comprises 11 adapter networks (including 2 projections), totaling 530MB (7% of base model size). Each adapter is a small MLP (typically 2–3 layers with hidden dimension 256–512).
Hardware.
All experiments run on a single NVIDIA Jetson AGX Orin (64GB unified memory). Training each adapter takes 2–5 minutes; full system training completes in under 2 hours.
Evaluation.
We introduce PrometheusExtract-132, a benchmark of 132 test cases with 203 expected facts, spanning 10 categories from clean singular inputs to adversarial multi-subject sentences. We report accuracy at multiple levels (L1: perspective, L3: relation, L5: full fact, L6: retrieval) with 95% confidence intervals computed via Wilson score.
5.2 Overall Results
Table 3 shows overall accuracy on PrometheusExtract-132.
Table 3: Overall accuracy on PrometheusExtract-132 (n=132 cases, 203 facts).
Level Metric Exact Partial
Extraction (per-fact)
L1 Perspective 97.7%
L2 Entity 82.3% 82.3%
L3 Relation 47.3%
L4 Value 80.8% 86.7%
L5 Full Fact 37.4% 39.9%
Retrieval (per-case)
L6 Query Answer 57.6% 59.6%
Retrieval Rate 69.5%
Key finding: Perspective detection (L1: 97.7%) and entity extraction (L2: 82.3%) are strong. Relation classification (L3: 47.3%) remains the primary bottleneck, responsible for 64 of 72 extraction errors (89%).
5.3 Results by Difficulty
The overall L6 accuracy of 57.6% (Table 3) masks a sharp performance split by input difficulty (Table 4).
Table 4: Accuracy by difficulty level with 95% confidence intervals.
Difficulty N L1 L5 L6 95% CI (L6)
Easy 54 100% 85.2% 94.4% [84.9%, 98.1%]
Medium 42 95.2% 28.6% 45.2% [31.2%, 60.1%]
Hard 36 97.2% 5.6% 19.4% [9.8%, 35.0%]
The system performs well on simple, clean inputs (94.4% retrieval) but degrades significantly on complex inputs. This is expected: the adapter pipeline was designed for clear factual statements, not adversarial or ambiguous text.
5.4 Results by Category
Table 5 breaks down performance across input categories.
Table 5: Accuracy by input category (L6 = end-to-end retrieval).
Category N L1 L5 L6
singular_clean 32 100% 93.8% 100%
clean 20 100% 80.0% 85.0%
singular_messy 16 100% 50.0% 68.8%
messy 16 93.8% 18.8% 31.2%
multi_temporal 8 100% 37.5% 50.0%
multi_domain 8 100% 0% 37.5%
multi_perspective 8 100% 0% 37.5%
multi_sentence 8 100% 0% 25.0%
multi_object 8 100% 0% 0%
multi_subject 8 75% 0% 0%
Strong: Perspective detection is now 97.7% overall (was 83.3% before UserAdapter fix). singular_clean achieves 100% L6 retrieval. Failure modes: multi_object (0%), multi_subject (0%)—the system cannot handle multiple subjects or objects in one sentence.
5.5 Ablation Studies
5.5.1 Stage-Wise vs End-to-End Training
In early development, we attempted end-to-end training of all 11 adapters jointly. This consistently collapsed to near-random performance (roughly 1/11 accuracy for 11 relation types). We attribute this to:
  • Credit assignment: Errors at extraction propagate through storage and retrieval, making gradient signals uninformative.
  • Gradient interference: Updates benefiting one adapter can degrade another’s performance.
Stage-wise training (each adapter on its own proxy task) avoids these issues. Note: We present this as an observation from development rather than a controlled experiment, as end-to-end training was abandoned early.
5.5.2 Identity V vs Learned V Encoder
We compare using lm_head.weight directly against a learned MLP:
V Encoder Train (n=68) Test (n=8)
Learned MLP 100% 0% (0/8)
Identity V 87.5% (7/8)
The learned V encoder memorizes training examples but fails completely on held-out facts. Identity V, requiring no training, generalizes because it uses the model’s native vocabulary representations. Limitation: n=8 test samples is insufficient for statistical confidence; this is a qualitative observation supporting the architectural choice.
5.5.3 With vs Without Projection
Hidden state collapse (Section 4.5) motivates learned projections. On a set of 500 word pairs, raw hidden states show 0.98+ cosine similarity for semantically distinct words. After projection, similarity drops to 0.09, matching the structure of embed_tokens.
5.6 Qualitative Examples
Table 6 shows example interactions demonstrating memory extraction, multi-hop reasoning, and generation.
Table 6: Example conversation showing extraction, direct queries, and multi-hop reasoning. Memory operations shown in italics.
Turn User Input / System Response Memory Operation
[…]
Response You work as an architect. Inject via Identity V
[…]
Response Mark is a doctor. Inject: “doctor”
Multi-hop Reasoning Trace.
For the query “Where does my husband work?”, the system performs:
  1. 1.
    Hop 0: IntermediateRelationDetector identifies “spouse” relation from “my husband”
  2. 2.
    Lookup: user $\rightarrow$ spouse $\rightarrow$ Mark
  3. 3.
    HopController: CONTINUE (Mark is an entity)
  4. 4.
    Hop 1: IntermediateRelationDetector identifies “city” from “where…work”
  5. 5.
    Lookup: Mark $\rightarrow$ city $\rightarrow$ NOT_FOUND
  6. 6.
    AnswerSelector: Return partial answer with profession (doctor) as fallback
6 Discussion
6.1 Why Stage-Wise Training Succeeds
The success of stage-wise training warrants explanation. We hypothesize three factors:
Direct supervision.
Each adapter receives clear, task-specific gradients. The BetaGate learns “is this an entity?” from direct labels, not from how its output eventually affects generation ten components downstream.
No interference.
When training one adapter, others are frozen. This prevents the oscillation that occurs when improving component A degrades component B, triggering updates that degrade A, and so on.
Compositionality.
Simple components compose into complex behavior. A BetaGate that perfectly detects entities and a RelationClassifier that perfectly classifies relations will together perfectly extract entity-relation pairs—no joint training required.
This suggests a broader principle: memory systems may be better viewed as compositions of specialists than as monolithic networks. Each specialist masters one skill; the system architect (here, Python code) handles composition.
6.2 The Identity V Insight
The success of using lm_head.weight as memory values reveals something fundamental about transformer language models. During pretraining, the model learned to map hidden states to vocabulary probabilities. Row $i$ of lm_head is precisely the direction that, when projected onto by a hidden state, produces high probability for token $i$.
This means lm_head already encodes “what direction produces each word”—exactly what we need for memory injection. By using these rows directly, we inherit the model’s own understanding of its vocabulary. Any token the model can output, we can inject.
[…]
6.3 Hidden State Geometry and Collapse
Our discovery that hidden states collapse semantically distinct concepts has implications beyond memory systems. Any application that compares transformer hidden states via cosine similarity—retrieval, clustering, classification—may suffer from this effect.
The 86%/14% split between shared and distinctive variance suggests transformers learn strong abstractions (“relationship noun”) that dominate the representation. This is likely beneficial for generation—the model can generalize across instances of a category—but problematic for discrimination.
The fix is straightforward: learn a projection that recovers distinction, using embed_tokens (which preserves distinction) as supervision. This technique may benefit other applications of hidden state comparison.
6.4 Limitations
We emphasize the limitations of our work:
  1. 1.
    Single base model: All experiments use Qwen3-4B. We have not validated transfer to other model families.
  2. 2.
    No benchmark comparisons: We do not evaluate on standard memory benchmarks such as LoCoMo or BABILong.
  3. 3.
    Partially learned routing: While all decision boundaries are learned, the composition logic (control flow) remains in Python.
  4. 4.
    Scale: Testing limited to 88 entities. Scaling behavior unknown.
  5. 5.
    Injection layer sensitivity: The layer 35 (97% depth) injection point was empirically tuned; theoretical understanding is lacking.
  6. 6.
    Single-token injection: Current implementation injects one token at a time. Multi-token answers (e.g., “New York”) require extension.
6.5 What We Are Not Claiming
We do not claim that Prometheus Mind outperforms existing memory systems on standard benchmarks, nor that our adapters transfer without modification to other model families. Our contribution is demonstrating that memory can be retrofitted to frozen models, and identifying principles that make this work.
7 Conclusion
We presented Prometheus Mind, demonstrating that memory can be retrofitted to a frozen language model without architectural changes—achieving 94.4% retrieval accuracy on clean inputs using only 11 modular adapters (7% parameter overhead) that are fully reversible.
The journey to this result required solving four problems, each yielding a reusable insight:
  1. 1.
    Extraction without labels: Contrastive Direction Discovery finds semantic directions by contrasting minimal pairs without supervision.
  2. 2.
    Reliable training: Stage-wise training of each adapter on simple proxy tasks succeeds where end-to-end training collapses.
  3. 3.
    Zero-training injection: lm_head.weight rows are already optimal value vectors—learned encoders fail to generalize.
  4. 4.
    Hidden state recovery: Transformer hidden states collapse distinct concepts, but projections trained on embed_tokens can recover distinction.
Broader Implications.
While demonstrated for memory, these principles may generalize to any capability expressible as routing to existing representations. The frozen model contains latent capabilities; adapters provide access patterns. This suggests a paradigm of modular cognition: composable, reversible capability modules over a fixed base. The value of reversibility extends beyond convenience—it enables A/B testing, rollback, and compliance with deployment constraints that prohibit weight modification.
Future Work.
Key directions include validating on other model families, evaluation on standard memory benchmarks (LoCoMo, BABILong), and investigating the theoretical basis for Identity V and injection layer selection.
Acknowledgments
This work represents human-AI collaboration. Mark Wind provided research direction, architectural decisions, and the core insight that memory should modify hidden states rather than input tokens. Claude (Anthropic) contributed implementation, systematic experimentation, writing assistance, and documentation. All code, experiments, and paper drafting involved iterative collaboration between human and AI.
The research was conducted entirely on consumer hardware (NVIDIA Jetson AGX Orin, 64GB), demonstrating that memory research is accessible without datacenter resources.
Appendix A Adapter Details
Table 7 provides complete details for all 11 adapters used in Prometheus Mind.
Table 7: Complete adapter inventory with architectures and training details.
Adapter Function Architecture Input N Acc
Extraction Adapters
[…]
IdentityAdapter v4 Self-intro detection MLP(256,2) Sentence emb 80 100%
UserAdapter v1 First-person detection MLP(256,2) Sentence emb 120 98%
FirstPersonValueMask v2 Multi-value extraction MLP(512,1) Per-token emb 150 100%
PairRelationClassifier v3 Relation (17 classes) Gated MLP(512,17) Value+context 200 100%
LocalWindowClassifier v1 Context-window relation MLP(256,16) $\pm$3 tokens 180 82%
Reasoning Adapters
IntermRelationDetector v4 Query relation (17) MLP(512,17) Projected sent 916 99.7%
HopController v3 Continue/stop MLP(256,2) Query+fact emb 50 100%
AnswerSelector v1 Entity vs value MLP(128,2) Query emb 60 100%
QueryEntityExtractor v1 Entity from question Attn+Linear Token embs 80 100%
Routing Adapters
TypeClassifier v1 Entity type (4) MLP(256,4) Entity emb 100 100%
RoutingAdapter v2 Slot assignment MLP(256,slots) Entity+rel 150 79%
MessageTypeClassifier v1 Q/statement/chat MLP(256,3) Sentence emb 100 95%
EntityClassifier v1 Is value entity? MLP(128,2) Value emb 100 100%
Projections (input dim 2560)
UniversalProjection v1 Fix word collapse MLP(2560,2560) Hidden state 500
SentenceProjection v1 Fix sentence collapse MLP(2560,2560) Sentence HS 683
Architecture Notation.
MLP($d_{\text{in}},d_{\text{hidden}},d_{\text{out}}$) denotes a two-layer network with ReLU activation: Linear($d_{\text{in}}\rightarrow d_{\text{hidden}}$) $\rightarrow$ ReLU $\rightarrow$ Linear($d_{\text{hidden}}\rightarrow d_{\text{out}}$). “Gated MLP” adds a learned gate combining value-primary and context features.
Training Details.
All adapters use AdamW optimizer with learning rate $10^{-4}$ and batch size 16. Training typically converges in 50–200 epochs (2–5 minutes on Jetson AGX Orin). Cross-entropy loss for classification; MSE loss for projections.
Total Parameters.
The 11 adapters total approximately 265M parameters (530MB at fp16), representing 7% of the 8GB base model.
Appendix B CDD Implementation Details
B.1 Discovered Attention Heads
Table 8 lists the attention heads discovered to be selective for each semantic role in Qwen3-4B.
Table 8: Attention heads selective for each semantic role, discovered via differential incoming attention analysis.
Role Heads (Layer, Head) Selection Criterion
Verb L6H2, L6H29, L5H20, L4H27, L5H23 $\mathbb{E}[\text{attn}_{\rightarrow\text{verb}}]>\mathbb{E}[\text{attn}_{\rightarrow\text{non-verb}}]$
Object L1H6, L6H21, L0H27, L5H18, L11H24 $\mathbb{E}[\text{attn}_{\rightarrow\text{obj}}]>\mathbb{E}[\text{attn}_{\rightarrow\text{noise}}]$
Temporal L1H1, L6H4, L0H21, L0H12, L6H11 $\mathbb{E}[\text{attn}_{\rightarrow\text{temp}}]>\mathbb{E}[\text{attn}_{\rightarrow\text{non-temp}}]$
Subject L0H3, L4H24, L0H30, L0H11, L0H23 Used for interjection filtering
B.2 Minimal Pair Examples
The following minimal pairs were used to compute CDD directions:
Subject-Object Direction.
  • “Alice saw Bob” vs “Carol saw Bob” (isolates subject)
  • “Alice saw Bob” vs “Alice saw Carol” (isolates object)
  • Difference: $\mathbf{d}_{\text{SO}}=\mathbb{E}[\mathbf{h}_{\text{subj}}-\mathbf{h}_{\text{verb}}]-\mathbb{E}[\mathbf{h}_{\text{obj}}-\mathbf{h}_{\text{verb}}]$
Head-Modifier Direction.
  • “The cat” $\rightarrow$ contrast “cat” vs “The”
  • “A doctor” $\rightarrow$ contrast “doctor” vs “A”
  • Difference: $\mathbf{d}_{\text{HM}}=\mathbb{E}[\mathbf{h}_{\text{noun}}]-\mathbb{E}[\mathbf{h}_{\text{det}}]$
B.3 Direction Computation Algorithm
1:Input: Set of minimal pairs $\{(s_{1}^{+},s_{1}^{-}),\ldots,(s_{n}^{+},s_{n}^{-})\}$
2:Output: Direction vector $\mathbf{d}$
3:for each pair $(s^{+},s^{-})$ do
4:  $\mathbf{h}^{+}\leftarrow$ GetHiddenState($s^{+}$, target_token)
5:  $\mathbf{h}^{-}\leftarrow$ GetHiddenState($s^{-}$, target_token)
6:  $\mathbf{\delta}_{i}\leftarrow\mathbf{h}^{+}-\mathbf{h}^{-}$
7:end for
[…]
10:return $\mathbf{d}$
B.4 Attention Head Discovery Algorithm
1:Input: Corpus with role annotations, model with $L$ layers and $H$ heads
2:Output: Set of role-selective heads
3:for layer $l=0$ to L1L-1 do
4:  for head $h=0$ to H1H-1 do
5:   $\text{role\_attn}\leftarrow[]$, $\text{other\_attn}\leftarrow[]$
6:   for each sentence in corpus do
7:    for each token position $i$ do
8:     $a_{i}\leftarrow\sum_{j>i}\text{Attention}[l,h,j,i]$ {Incoming attention}
9:     if token $i$ has target role then
10:      Append $a_{i}$ to role_attn
11:     else
12:      Append $a_{i}$ to other_attn
13:     end if
14:    end for
15:   end for
16:   $\Delta_{l,h}\leftarrow\text{mean}(\text{role\_attn})-\text{mean}(\text{other\_attn})$
17:  end for
18:end for
19:return Top-$k$ heads by $\Delta_{l,h}$
Appendix C Additional Experiments
C.1 Per-Relation Retrieval Accuracy
Table 9 shows retrieval accuracy broken down by relation type.
Table 9: Retrieval accuracy by relation type.
Relation Direct Multi-hop
identity 100% (5/5)
profession 100% (4/4)
city 86% (6/7) 100% (2/2)
spouse 100% (2/2) 100% (2/2)
sibling 100% (2/2) 75% (3/4)
works_with 100% (3/3) 100% (2/2)
friend 100% (2/2) 100% (1/1)
interest 75% (3/4)
belongs_to 100% (2/2)
C.2 Projection Training Convergence
Both projections converge rapidly when trained to match target similarity structures:
Table 10: Projection training convergence.
Projection Epochs Final Loss Similarity Match
UniversalProjection 150 0.008 $r=0.94$
SentenceProjection 200 0.012 $r=0.91$
C.3 Failure Case Analysis
We analyze failure cases to understand system limitations:
Third-Person Possessive Chains.
Query: “What does Alice’s colleague do?” (Alice $\rightarrow$ works_with $\rightarrow$ Bob $\rightarrow$ profession)
Failure mode: QueryEntityExtractor extracts “Alice’s colleague” as entity instead of “Alice”.
Root cause: Possessive constructions not in training data.
Ambiguous Relations.
Input: “I love hiking and my sister Sarah”
Failure mode: “Sarah” classified as interest instead of sibling.
Root cause: LocalWindowClassifier sees “love…Sarah” without “sister” in $\pm$3 window.
Interest Extraction.
Input: “I’m really into photography”
Failure mode: “photography” not extracted (FirstPersonValueMask misses it).
Root cause: Colloquial phrasing “into X” not in training distribution.
C.4 Injection Depth Sensitivity
We evaluated injection at different layer depths using two methods:
Hidden state blending (early experiments).
Direct addition of memory embeddings to hidden states showed optimal performance at 86% depth (layer 30/36), with degradation at both earlier and later layers. However, this method achieved only 91% accuracy even at optimal depth.
Attention K-V injection (final method).
The breakthrough came from injecting memory as key-value pairs into the attention mechanism at layer 35 (97% depth). Combined with first-token-only injection and Identity V, this achieves near-perfect retrieval accuracy. The key insight: at 97% depth, the model has built full context, and attention-based injection avoids the repetition issues that plagued direct hidden state modification at late layers.
Table 11: Accuracy by injection method and depth.
Method Depth Layer Accuracy
Hidden state blend 86% 30/36 91%
Hidden state blend 97% 35/36 62% (repetition)
Attention K-V 97% 35/36 $\sim$100%
The difference illustrates that how memory is injected matters as much as where. Attention-based injection at late layers succeeds where direct hidden state modification fails.
C.5 Memory Scale Testing
We tested with varying numbers of stored entities:
Table 12: Retrieval accuracy by memory size.
Entities Direct Accuracy Multi-hop Accuracy
10 100% 100%
25 96% 92%
50 94% 88%
88 91% 85%
Performance degrades gracefully with scale, though we have not tested beyond 88 entities. The primary bottleneck is the RoutingAdapter (79% accuracy), which occasionally assigns entities to incorrect memory slots